I've been having trouble understanding how to solve these, please help?
Given f(x) = x − 1/x and g(x) = 1/x, what is f(g(1/2))
Thank you

Answer by fractalier(6550)

f(g(1/2)) = f(2) = 2 - 1/2 = 3/2
since g(1/2) = 1/(1/2) = 2